OpenStudy (anonymous):

Hi, I am looking to learn Linear equations with 2 variables, I do not understand the concept. Can someone please help? Thanks

OpenStudy (shayaan_mustafa):

Hi antjojo2009 :) How are you? :)

OpenStudy (shayaan_mustafa):

Which concept regarding with this equation is not clear to you dear?

OpenStudy (anonymous):

Hi good

OpenStudy (anonymous):

I am not understanding how to work 2 variable linear equations

OpenStudy (shayaan_mustafa):

I am going for dinner. Until I come back, if you wait then I will help you thoroughly or if you want then you can bump it. Please let me go now then I will be back.

OpenStudy (shayaan_mustafa):

Until post your all problem.

OpenStudy (saifoo.khan):

MAy i help? @Shayaan_Mustafa

OpenStudy (shayaan_mustafa):

Yes saifoo.khan. Please. thanks

OpenStudy (shayaan_mustafa):

saifoo.khan will help you thoroughly. Good luck guys. Thanks once again. tc

OpenStudy (saifoo.khan):

@antjojo2009 are u still here?

OpenStudy (saifoo.khan):

Let me write things out. @antjojo2009

OpenStudy (anonymous):

Hi any help would be great

OpenStudy (saifoo.khan):

Well when we have equations with 2 variables we are basically looking for the value of each. Let's say if we are given, 5 caps and 2 watches cost 13 dollars. And 1 cap and 2 watches cost 9 dollars.

OpenStudy (anonymous):

still here

OpenStudy (saifoo.khan):

Now we are going to make two equations, let's suppose caps as x and watches as y. we can rewrite as, 5x + 2y = 13 -equation 1 1x + 2y = 9 -equation 2.

OpenStudy (saifoo.khan):

Now we can solve both equations to get the value of x and y. therefore, giving us the value of each cap and each watch. Do you understand?

OpenStudy (anonymous):

yes, i understand. The problem I am having is coming up with the answer

OpenStudy (saifoo.khan):

Ok, you may post your problem. i will explain it.

OpenStudy (anonymous):

y=5x+1 5y+4=7x

OpenStudy (saifoo.khan):

Do you know about substitution method?

OpenStudy (anonymous):

not really

OpenStudy (saifoo.khan):

umm, okay. in substituion method, we take one equation and make it the subject and then insert the value into another equation to solve for it, let me show you.

OpenStudy (saifoo.khan):

y=5x+1 -equation 1 5y+4=7x -equation 2 Equation 1 has "y" as subject , so we can ignore the first step. in step 2, we will insert the value of y in another equation (equation2) 5y + 4 = 7x Now replace y with the value of y in equation 1. 5(5x+1) + 4 = 7x. Do you copy till here?

OpenStudy (anonymous):

oh okay, yes

OpenStudy (saifoo.khan):

Cool, now solve the equation: 5(5x+1) + 4 = 7x. 25x + 5 + 4 = 7x after solving we willl get, x = -1/2 Got it?

OpenStudy (anonymous):

ok thanks a bunch

OpenStudy (saifoo.khan):

Hey! there's 1 more step left.

OpenStudy (saifoo.khan):

Now after you have the value of x. insert that value into any equation you like to get the value of y. Let me insert it into equation 1, y=5x+1 y = 5(-1/2) + 1 y = -5/2 + 1 y = -3/2

OpenStudy (anonymous):

ok